OpenStudy (anonymous):

Simplify (tanx + cosx + sinxtanx)/(secx + tanx) [answer: 1] Could someone please explain to me how to get the answer?

OpenStudy (anonymous):

convert all identies in sinx and cosx and then solve.i hope u will find ans urself

OpenStudy (jdoe0001):

girlnotonfire do you happen to have your trig identities sheet?

OpenStudy (anonymous):

Just the ones in my math book.

OpenStudy (jdoe0001):

ok.... so you know what hmm say \(\bf cos^2(x)+sin^2(x)=?\) is right?

OpenStudy (anonymous):

1

OpenStudy (jdoe0001):

ok so as WHassan suggested, let's expand the identities to just cosine and sine identities... so

OpenStudy (jdoe0001):

one sec

OpenStudy (anonymous):

u should not give direct ans

OpenStudy (jdoe0001):

\(\bf \color{blue}{ tan(x)=\frac{sin(x)}{cos(x)}\qquad \qquad sec(x)=\frac{1}{cos(x)}\qquad cos^2(x)+sin^2(x)=1}\\ \quad \\ \quad \\ \cfrac{tan(x) + cos(x) + sin(x)tan(x)}{sec(x) + tan(x)}\implies \cfrac{{\color{blue}{ \frac{sin(x)}{cos(x)}}}+cos(x)+sin(x)\frac{sin(x)}{cos(x)}}{{\color{blue}{ \frac{1}{cos(x)}}}+\frac{sin(x)}{cos(x)}}\\ \quad \\ \implies \cfrac{ \frac{sin(x)}{cos(x)}+cos(x)+\frac{sin(x)^2}{cos(x)} }{\frac{1}{cos(x)}+\frac{sin(x)}{cos(x)}}\implies \cfrac{\frac{sin(x)+cos^2(x)+sin^2(x)}{cos(x)}}{\frac{1+sin(x)}{cos(x)}} \)

OpenStudy (jdoe0001):

\(\bf {\color{blue}{ tan(x)=\frac{sin(x)}{cos(x)}}\qquad \qquad sec(x)=\frac{1}{cos(x)}\qquad cos^2(x)+sin^2(x)=1}\\ \quad \\ \quad \\ \cfrac{ \frac{sin(x)}{cos(x)}+cos(x)+\frac{sin(x)^2}{cos(x)} }{\frac{1}{cos(x)}+\frac{sin(x)}{cos(x)}}\implies \cfrac{\frac{sin(x)+{\color{blue}{ cos^2(x)+sin^2(x)}}}{cos(x)}}{\frac{1+sin(x)}{cos(x)}}\) so... notice that.... and replace by 1, then do the fractions, keeping in mind that \(\bf \cfrac{\quad \frac{a}{b}\quad }{\frac{c}{d}}\implies \cfrac{a}{b}\cdot \cfrac{d}{c}\)

OpenStudy (anonymous):

Oh okay, that makes since. It simplifies to \[\cos^2x+\sin^2x\] so it equals 1.

OpenStudy (anonymous):

sense*

OpenStudy (jdoe0001):

\(\bf {\color{blue}{ tan(x)=\frac{sin(x)}{cos(x)}\qquad \qquad sec(x)=\frac{1}{cos(x)}\qquad cos^2(x)+sin^2(x)=1}}\\ \quad \\ \quad \\ \cfrac{ \frac{sin(x)}{cos(x)}+cos(x)+\frac{sin(x)^2}{cos(x)} }{\frac{1}{cos(x)}+\frac{sin(x)}{cos(x)}}\implies \cfrac{\frac{sin(x)+{\color{blue}{ cos^2(x)+sin^2(x)}}}{cos(x)}}{\frac{1+sin(x)}{cos(x)}} \)

OpenStudy (anonymous):

Thank you very much for your help!!

OpenStudy (jdoe0001):

right... so .. \(\bf {\color{blue}{ tan(x)=\frac{sin(x)}{cos(x)}\qquad \qquad sec(x)=\frac{1}{cos(x)}\qquad cos^2(x)+sin^2(x)=1}}\\ \quad \\ \quad \\ \cfrac{\frac{sin(x)+{\color{blue}{ cos^2(x)+sin^2(x)}}}{cos(x)}}{\frac{1+sin(x)}{cos(x)}}\implies \cfrac{\frac{sin(x)+{\color{blue}{ 1}}}{cos(x)}}{\frac{1+sin(x)}{cos(x)}}\\ \quad \\ \implies \cfrac{sin(x)+{\color{blue}{ cos^2(x)+sin^2(x)}}}{\cancel{cos(x)}}\cdot \cfrac{\cancel{cos(x)}}{1+sin(x)}\)

OpenStudy (jdoe0001):

woopps... anyhow I got the wrong ... fractoin... but \(\bf \cfrac{sin(x)+{\color{blue}{ 1}}}{\cancel{cos(x)}}\cdot \cfrac{\cancel{cos(x)}}{1+sin(x)}\)